please help.find the roots of the quadratic equation y=x^2+7x-18 by factoring Thanks
Answer by oscargut(2103)

y=x^2+7x-18 =x^2-2x+9x-18=x(x-2)+9(x-2)=(x+9)(x-2)
then if (x+9)(x-2)=0 roots are x=-9 and x=2